Use slope intercept form to write the equation of the line: passing through the points (5,6) and (7,10)
Answer by nyc_function(2741)

First find the slope.
Let m = slope
m = (10 - 6)/(7 - 5)
m = 4/2
m = 2
The slope is 2.
We now use the point-slope formula. In the point-slope formula, we use one of the given points and the slope.
y - b = m(x - a)....This is the point-slope formula.
I will use (5,6), where a = 5 and b = 6. The slope is 2.
y - 6 = 2(x - 5)
y - 6 = 2x - 10
y = 2x - 10 + 6
y = 2x - 4
Did you follow?